# Effective Learning with Node Perturbation in Multi-Layer Neural Networks

### v5

@@ Appendix A Weights propagate noise correlations
Traditional NP updates have a number of sources of error. One reason is that use of the injected noise $\mathbf{\epsilon}_{l}$ to compute the learning signal ignores how the output activity of layer $l$ is not only impacted by the noise injected directly into it, but also by the cumulative effect of perturbations added to previous layers. If the noise is randomly sampled from a Gaussian distribution with mean zero, one might be tempted to simply assume that the perturbations from previous layers cancel out in expectation, but this assumption ignores the correlations introduced into these perturbations by the network’s weight matrices at all preceding layers. Multiplying a random vector by a non-orthogonal matrix will introduce a non-random covariance structure into it. To see why, consider a transformation $\mathbf{W}\mathbf{x}=\mathbf{y}$, where $\mathbf{W}$ is a randomly initialized weight matrix and $\mathbf{x}$ is an uncorrelated input vector for which it holds that $\langle\mathbf{x}\mathbf{x}^{\top}\rangle=\mathbf{I}$ with $\mathbf{I}$ the identity matrix. The covariance matrix of $\mathbf{y}$ can be described as
@@ Appendix C Derivation of activity-based node perturbation
Appendix D Bias in NP compared to ANP
In this section, we analyze the deviation of NP’s and ANP’s updates from the true gradient in a simple two-layer neural network, focusing on the effects of noise propagation through the network’s layers. Unlike existing work in the literature, we do not assume infinitely small perturbations, or a linear network, to provide a quantification of bias in NP’s and ANP’s updates. In practice, perturbations applied to networks cannot be infinitely small and are even limited by precision of floating point representations to be larger than some minimum size.
@@ Appendix D Bias in NP compared to ANP
$\mathbf{y}_{2}=\mathbf{W}_{2}f(\mathbf{W}_{1}\mathbf{x})\,.$
@@
$\tilde{\mathbf{y}}_{2}=\mathbf{W}_{2}\big{(}f(\mathbf{W}_{1}\mathbf{x}+\mathbf{v}_{1})+\mathbf{v}_{2}\big{)}\,.$
@@
Given an arbitrary loss function $L$, the loss differential is given by $L(\tilde{\mathbf{y}}_{2})-L(\mathbf{y}_{2})$. NP’s updates are based on a product between this loss differential and the noise. Examining this for matrix $\mathbf{W}_{2}$, this corresponds to a product between the noise differential and vector $\mathbf{v}_{2}$. To fully show the impact of this update, we expand the loss differential by a Taylor expansion beyond the first term (i.e. assuming that our loss difference is not infinitely small or linearizable) and then further expand the activity difference by Taylor expansion. As such,
@@
$\langle\Delta\mathbf{W}_{2}^{\text{NP}}\rangle$ $=\frac{1}{h}\left\langle(L(\tilde{\mathbf{y}}_{2})-L(\mathbf{y}_{2}))\cdot\mathbf{v}_{2}\right\rangle f(\mathbf{y}_{1})^{\top}$
$=\frac{1}{h}\Bigg{\langle}\mathbf{v}_{2}\Bigg{[}\nabla_{\mathbf{y}_{2}}L^{\top}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})+\frac{1}{2!}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\nabla^{2}_{\mathbf{y}_{2}}L\;(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})+\frac{1}{3!}\cdots\Bigg{]}^{\top}\Bigg{\rangle}f(\mathbf{y}_{1})^{\top}$
@@
$=\frac{1}{h}\left\langle\mathbf{v}_{2}\left[\nabla_{\mathbf{y}_{2}}L^{\top}\left(\mathbf{v}_{2}+\nabla_{\mathbf{y}_{1}}\mathbf{y}_{2}^{\top}\mathbf{v}_{1}+\frac{1}{2!}\mathbf{v}_{1}^{\top}\nabla^{2}_{\mathbf{y}_{1}}\mathbf{y}_{2}\mathbf{v}_{1}+\frac{1}{3!}\cdots\right)\cdots\right]^{\top}\right\rangle f(\mathbf{y}_{1})^{\top}$
@@
$=\frac{1}{h}\left\langle\mathbf{v}_{2}\left[\left(\mathbf{v}_{2}+\nabla_{\mathbf{y}_{1}}\mathbf{y}_{2}^{\top}\mathbf{v}_{1}+\frac{1}{2!}\mathbf{v}_{1}^{\top}\nabla^{2}_{\mathbf{y}_{1}}\mathbf{y}_{2}\mathbf{v}_{1}+\frac{1}{3!}\cdots\right)^{\top}\nabla_{\mathbf{y}_{2}}L\cdots\right]\right\rangle f(\mathbf{y}_{1})^{\top}$
$=\frac{1}{h}\bigg{\langle}\mathbf{v}_{2}\mathbf{v}_{2}^{\top}+\mathbf{v}_{2}\mathbf{v}_{1}^{\top}\nabla_{\mathbf{y}_{1}}\mathbf{y}_{2}+\frac{1}{2!}\mathbf{v}_{2}\mathbf{v}_{1}^{\top}\nabla^{2}_{\mathbf{y}_{1}}\mathbf{y}_{2}^{\top}\mathbf{v}_{1}+\frac{1}{3!}\cdots\bigg{\rangle}\nabla_{\mathbf{y}_{2}}L\;f(\mathbf{y}_{1})^{\top}$
=$\cdots$= ⋯
The above Taylor expansion demonstrates that above the linear terms ($\mathbf{v}_{2}\mathbf{v}_{2}^{\top}+\mathbf{v}_{2}\mathbf{v}_{1}^{\top}\nabla_{\mathbf{y}_{1}}\mathbf{y}_{2}\approx\mathbf{I}+\mathbf{0}$, in the ideal noise case), there are terms involving the correlation between noise vector $\mathbf{v}_{2}$ and higher order terms invoving $\mathbf{v}_{1}$, beginning with the Hessian but also including third, fourth and higher derivative terms. We cannot, in general, say anything about the unbiasedness of these terms and these are sources of unexplained error in existing work.
By comparison, ANP’s Taylor expansion has a much more interpretable correlation structure. Again looking at weight updates for layer two:
$\langle\Delta\mathbf{W}_{2}^{\text{ANP}}\rangle$ $=\frac{1}{h}\Big{\langle}\big{(}L(\tilde{\mathbf{y}}_{2})-L(\mathbf{y})\big{)}\cdot(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})\Big{\rangle}f(\mathbf{y}_{1})^{\top}$
@@
$=\frac{1}{h}\Big{\langle}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})\Big{[}\nabla_{\mathbf{y}_{2}}L^{\top}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})+\frac{1}{2!}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\nabla^{2}_{\mathbf{y}_{2}}L(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})+\frac{1}{3!}\dots\Big{]}^{\top}\Big{\rangle}f(\mathbf{y}_{1})^{\top}$
@@
$=\frac{1}{h}\Big{\langle}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\nabla_{\mathbf{y}_{2}}L+\frac{1}{2!}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\nabla^{2}_{\mathbf{y}_{2}}L^{\top}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})+\frac{1}{3!}\dots\Big{]}^{\top}\Big{\rangle}f(\mathbf{y}_{1})^{\top}\,.$
This simplifies all assumptions regarding higher-order noise propagation terms, and reduces the algorithms bias down to the correlation in overall activity differences ($\langle(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\rangle$). This is a much more interpretable bias term which could even be targeted directly to be made diagonal if a practitioner was interested in doing so.
@@ Appendix I Peak accuracies
Figure 8 shows train and test loss for (D)NP, (D)ANP and (D)BP for the SARCOS dataset (Vijayakumar & Schaal, 2000). The task is to solve inverse dynamics problem for a seven degrees-of-freedom SARCOS anthropomorphic robot arm by mapping from a 21-dimensional input space (7 joint positions, 7 joint velocities, 7 joint accelerations) to the corresponding 7 joint torques. The network is a fully connected architecture with two 32-unit hidden layers and was trained with MSE loss and the leaky ReLU activation function. The network was trained for 500 epochs and results were averaged over three random seeds. A learning rate search for each algorithm was started at $10^{-7}$, doubling the learning rate until the algorithm became unstable. A decorrelation learning rate of $10^{-4}$ was used for DNP and DANP and $10^{-3}$ for DBP.

### v6

@@ Appendix A Weights propagate noise correlations
Traditional NP updates have a number of sources of error. One reason is that use of the injected noise $\mathbf{\epsilon}_{l}$ to compute the learning signal ignores how the output activity of layer $l$ is not only impacted by the noise injected directly into it, but also by the cumulative effect of perturbations added to previous layers. If the noise is randomly sampled from a Gaussian distribution with mean zero, one might be tempted to simply assume that the perturbations from previous layers cancel out in expectation, but this assumption ignores the correlations introduced into these perturbations by the network’s weight matrices at all preceding layers. Multiplying a random vector by a non-orthogonal matrix will introduce a non-random covariance structure into it.
To see why, consider a transformation $\mathbf{W}\mathbf{x}=\mathbf{y}$, where $\mathbf{W}$ is a randomly initialized weight matrix and $\mathbf{x}$ is an uncorrelated input vector for which it holds that $\langle\mathbf{x}\mathbf{x}^{\top}\rangle=\mathbf{I}$ with $\mathbf{I}$ the identity matrix. The covariance matrix of $\mathbf{y}$ can be described as
@@ Appendix C Derivation of activity-based node perturbation
Appendix D Gradient Alignment of NP vs ANP
In this section, we analyze the deviation of NP’s and ANP’s updates from the true gradient in a simple two-layer neural network, focusing on the effects of noise propagation through the network’s layers.
In practice, perturbations applied to networks cannot be infinitely small and are even limited by precision of floating point representations to be larger than some minimum size. Nonetheless, we here assume that perturbations are sufficiently small that networks can be assumed to have been linearized.
@@ Appendix D Gradient Alignment of NP vs ANP
$\mathbf{y}_{2}=\mathbf{W}_{2}f(\mathbf{y}_{1})=\mathbf{W}_{2}f(\mathbf{W}_{1}\mathbf{x})\,.$
@@
$\tilde{\mathbf{y}}_{2}=\mathbf{W}_{2}\big{(}f(\mathbf{y}_{1}+\mathbf{v}_{1})\big{)}+\mathbf{v}_{2}=\mathbf{W}_{2}\big{(}f(\mathbf{W}_{1}\mathbf{x}+\mathbf{v}_{1})\big{)}+\mathbf{v}_{2}\,.$
@@
Given an arbitrary loss function $L$, the loss differential is given by $L(\tilde{\mathbf{y}}_{2})-L(\mathbf{y}_{2})$. NP’s updates are based on a product between this loss differential and the noise. Examining this for matrix $\mathbf{W}_{2}$, this corresponds to a product between the noise differential and vector $\mathbf{v}_{2}$. To fully show the impact of this update, we expand the loss differential by a first-order Taylor expansion.
As such,
@@
$\langle\Delta\mathbf{W}_{2}^{\text{NP}}\rangle$ $=\frac{1}{\sigma^{2}}\left\langle\delta L\cdot\mathbf{v}_{2}\right\rangle f(\mathbf{y}_{1})^{\top}$
@@
$=\frac{1}{\sigma^{2}}\left\langle\mathbf{v}_{2}\,(L(\tilde{\mathbf{y}}_{2})-L(\mathbf{y}_{2}))^{\top}\right\rangle f(\mathbf{y}_{1})^{\top}$
@@
$\approx\frac{1}{\sigma^{2}}\left\langle\mathbf{v}_{2}\,(\nabla_{\mathbf{y}_{2}}L^{\top}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2}))^{\top}\right\rangle f(\mathbf{y}_{1})^{\top}$
@@
$\approx\frac{1}{\sigma^{2}}\left\langle\mathbf{v}_{2}\,(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\nabla_{\mathbf{y}_{2}}L\right\rangle f(\mathbf{y}_{1})^{\top}$
@@
$\approx\frac{1}{\sigma^{2}}\left\langle\mathbf{v}_{2}\,(\mathbf{v}_{2}+\nabla_{\mathbf{y_{1}}}\mathbf{y_{2}}\circ\mathbf{v}_{1})^{\top}\right\rangle\nabla_{\mathbf{y}_{2}}L\,f(\mathbf{y}_{1})^{\top}$
$\approx\frac{1}{\sigma^{2}}\left\langle\mathbf{v}_{2}\mathbf{v}_{2}^{\top}+\mathbf{v}_{2}\mathbf{v}_{1}^{\top}\circ\nabla_{\mathbf{y_{1}}}\mathbf{y_{2}}^{\top}\right\rangle\nabla_{\mathbf{y}_{2}}L\,f(\mathbf{y}_{1})^{\top}$
In regular node perturbation, it is thereafter supposed that the terms in the angular brackets cancel down to the identity, $\langle\mathbf{v}_{2}\mathbf{v}_{2}^{\top}\rangle=\mathbf{I}$ and $\langle\mathbf{v}_{2}\mathbf{v}_{1}^{\top}\rangle=\mathbf{0}$, given our initial assumptions about noise sampling. However, in the case in which only a single (or relatively few) noise samples are used to compute an NP update, both of the terms related to the correlation between nodes noise at the second layer, $\langle\mathbf{v}_{2}\mathbf{v}_{2}^{\top}\rangle$, and the correlation between node noise between the first and second layers, $\langle\mathbf{v}_{2}\mathbf{v}_{1}^{\top}\rangle$, contribute an error to this update. Most notably, any error in the first term contributes to the gradient descent update being multiplied by a positive semi-definite matrix, which still ensures that the loss will be reduced. However any error in the second term, i.e. $\langle\mathbf{v}_{2}\mathbf{v}_{1}^{\top}\rangle\neq\mathbf{0}$, is not positive semi-definite and can throw off the update of NP away from reducing the loss.
Carrying out the same analysis for the ANP rule,
$\langle\Delta\mathbf{W}_{2}^{\text{ANP}}\rangle$ $=\Big{\langle}\gamma\big{(}L(\tilde{\mathbf{y}}_{2})-L(\mathbf{y})\big{)}\cdot(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})\Big{\rangle}f(\mathbf{y}_{1})^{\top}$
$\approx\Big{\langle}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})\Big{[}\nabla_{\mathbf{y}_{2}}L^{\top}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})\Big{]}^{\top}\Big{\rangle}f(\mathbf{y}_{1})^{\top}$
$\approx\Big{\langle}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\nabla_{\mathbf{y}_{2}}L\Big{\rangle}f(\mathbf{y}_{1})^{\top}$
$\approx\Big{\langle}(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\Big{\rangle}\nabla_{\mathbf{y}_{2}}Lf(\mathbf{y}_{1})^{\top}$
where $\gamma$ represents the normalizing factor required for the update rule. ANP’s relationship to the true gradient is far simpler than that of NP, only depends upon the output activations difference, and reduces the algorithms bias down to the correlation in overall activity difference at this layer ($\langle(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})(\tilde{\mathbf{y}}_{2}-\mathbf{y}_{2})^{\top}\rangle$). This correlation will strictly contribute a positive semi-definite matrix multiplication against the gradient update (regardless of how many noise samples are used for the averaging), thus ensuring that the loss is nonetheless reduced. It is also a much more interpretable bias term which could even be targeted directly to be made diagonal if a practitioner was interested in sampling to ensure this property.
Note that the comparisons in this section lean upon on the well known property that positive semi-definite matrices never modify a vector to be more than $90^{\circ}$ away from it’s initial orientation.
@@ Appendix I Peak accuracies
Figure 7 shows train and test loss for (D)NP, (D)ANP and (D)BP for the SARCOS dataset [Vijayakumar and Schaal, 2000]. The task is to solve inverse dynamics problem for a seven degrees-of-freedom SARCOS anthropomorphic robot arm by mapping from a 21-dimensional input space (7 joint positions, 7 joint velocities, 7 joint accelerations) to the corresponding 7 joint torques. The network is a fully connected architecture with two 32-unit hidden layers and was trained with MSE loss and the leaky ReLU activation function. The network was trained for 500 epochs and results were averaged over three random seeds. A learning rate search for each algorithm was started at $10^{-7}$, doubling the learning rate until the algorithm became unstable. A decorrelation learning rate of $10^{-4}$ was used for DNP and DANP and $10^{-3}$ for DBP.